how many liters of a 35% antifreeze solution should be mixed with 12 liters of a 25% antifreeze solution to make a solution that is 26% antifreeze? Round to nearest hundreth.
Answer by stanbon(75887)

how many liters of a 35% antifreeze solution should be mixed with 12 liters of a 25% antifreeze solution to make a solution that is 26% antifreeze? Round to nearest hundreth.
------------------
Equation:
alcohol + alcohol = alcohol
0.35x + 0.25*12 = 0.26(x+12)
Multiply thru by 100 to get:
35x + 25*12 = 26x + 26*12
9x = 12
x = 4/3 liter (amt. of 35% solution needed in the mixture)
=========================
Cheers,
Stan H.